# On an approach to canonicalizing elliptic Feynman integrals

### v2

@@ Acknowledgements
We would like to thank Ming Lian for collaborations in the early stage of this work. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12375097, 12347103, Science Foundation of China University of Petroleum, Beijing (No.2462025YJRC019), and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Construct $\varepsilon$-factorized basis with constant period matrix
In Sec. 2.3, we apply the derivative basis method Pogel:2022vat to construct the canonical basis. However, it is not the only method to find an $\varepsilon$-factorized basis. With a pre-canonical basis whose connection matrix in a special linear form, we can also try the strategy in vonManteuffel:2017hms ; Frellesvig:2021hkr .
As already pointed out in Sec. 2.2, the general structure of $\bm{A}^{(0)}$ motivates us to do the following linear transformation
$\phi_{i-3}\to\big{(}\,\mathcal{T}_{1}\cdot\vec{\phi}\,\big{)}_{i-3}=\phi_{i-3}+\left[F(\sin^{-1}a_{i},\lambda)-E(\sin^{-1}a_{i},\lambda)\right]\phi_{1}-F(\sin^{-1}a_{i},\lambda)\phi_{n-1}\,,$ (105)
for $15\leq i\leq n+$ 1, with $a_{i}=\sqrt{e_{i}/\lambda}$, and we have implicitly introduced the rotation matrix $\mathcal{T}_{1}$ acting on the basis. The transformation carries over to the integral level, namely, $I_{i-3}\to\big{(}\,\mathcal{T}_{1}\cdot\vec{I}\,\big{)}_{i-3}$, and we will suppress it for brevity. After such a transformation, it is straightforward to verify that the derivatives (with respect to $\lambda$ or $e_{i}$) of master integrals $I_{j}$’s with $23\leq j\leq n-$ 2, are $\varepsilon$-factorized under the maximal cut.
Now, the remaining task is to further transform $\phi_{1}$ and $\phi_{n-1}$, to rotate the corresponding $2\times 2$ block of $\bm{A}^{(0)}$ away, which is only $\lambda$ relevant. To do that, we only need to consider the corresponding $2\times 2$ block of the period matrix for $\phi_{1}$ and $\phi_{n-1}$, which is
@@
${\cal T}={\cal T}_{2}\cdot{\cal T}_{1}=\begin{pmatrix}E(1-\lambda)&0&\cdots&0&-K(1-\lambda)\\ F(\sin^{-1}a_{5},\lambda)-E(\sin^{-1}a_{5},\lambda)&1&\cdots&0&-F(\sin^{-1}a_{5},\lambda)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ F(\sin^{-1}a_{\infty},\lambda)-E(\sin^{-1}a_{\infty},\lambda)&0&\cdots&1&-F(\sin^{-1}a_{\infty},\lambda)\\ E(\lambda)-K(\lambda)&0&\cdots&0&K(\lambda)\end{pmatrix}\,.$ (108)
We’ve briefly discussed the asymptotic behaviors of the pre-canonical basis in the degenerate limit $c_{1}\to c_{2}$ or $\lambda\to 1$. It’s natural to consider the asymptotic behaviors of the $\varepsilon$-factorized basis as well. It suffices to study the asymptotic behaviors of the elliptic integrals appearing in the transformation matrices ${\cal T}_{1}$ and ${\cal T}_{2}$. We have the following:
@@
Note that $K(\lambda)$ and $F(\sin^{-1}a_{i},\lambda)$ are singular as $c_{1}\to c_{2}$, and they are logarithmically singular in this limit. These singularities do not cause problems for the basis integrals, since they cancel due to the fact that $\phi_{1}=\phi_{2}$ when $c_{1}\to c_{2}$; see Eq. (53). This cancellation also implies that the first $\varepsilon$-factorized basis integral vanishes in this limit. Similar to the pre-canonical basis, other $\varepsilon$-factorized basis integrals are ${\rm d}\log$-forms asymptotically in the neighborhood of the degenerate points $c_{1}=c_{2}$, as desired. From the demonstration above, we can see the construction in Sec. 2.3 is essentially different from the procedure here from two related perspectives. On the one hand, we introduce $\varepsilon$-dependence in the fiber transformations in the former construction, while the construction here does not. On the other hand, there are only complete elliptic integrals with a unique modulus arguments in the former case, while here we introduce incomplete elliptic integrals and both of complementary moduli arguments are needed. As a consequence, the construction here will introduce more elliptic integrals and undesirable asymptotic properties for the transformation. More relevant discussions can also be found in Frellesvig:2023iwr .
@@ B.1.1 The construction of pre-canonical basis in the Jacobi form
$\phi_{1}$ $=\frac{{\rm d}t}{\sqrt{P_{J}(t)}}={\rm d}K(\sin^{-1}t,k^{2})\,,$ (114a)
$\phi_{i-3}$ $=\frac{1}{t-a_{i}}\frac{\sqrt{P_{J}(a_{i})}}{\sqrt{P_{J}(t)}}{\rm d}t=\frac{1}{2}{\rm d}\log\left(\frac{1+\sqrt{\frac{(1-t^{2})(1-k^{2}a_{i}^{2})}{(1-a_{i}^{2})(1-k^{2}t^{2})}}}{1-\sqrt{\frac{(1-t^{2})(1-k^{2}a_{i}^{2})}{(1-a_{i}^{2})(1-k^{2}t^{2})}}}\right)-\frac{\sqrt{P_{J}(a_{i})}}{a_{i}}{\rm d}\Pi(1/a_{i}^{2},\sin^{-1}t,k^{2})\,,$ (114b)
$\phi_{n-2}$ $=\frac{t}{\sqrt{P_{4J}(t)}}{\rm d}t=\frac{1}{2}{\rm d}\log\left(\frac{1+k\sqrt{\frac{1-t^{2}}{1-k^{2}t^{2}}}}{1-k\sqrt{\frac{1-t^{2}}{1-k^{2}t^{2}}}}\right)\,,$ (114c)
$\phi_{n-1}$ $=\frac{1-k^{2}t^{2}}{\sqrt{P_{J}(t)}}{\rm d}t={\rm d}E(\sin^{-1}t,k^{2})\,,$ (114d)
where $i=5,\cdots,n$ and the next to last basis differential $\phi_{n-2}$ is a ${\rm d}\log$-form. Then we can find that the differential equation is in a special linear form. The $\varepsilon^{0}$-part of the connection matrix with respect to $k$ is
@@
B.2 The canonical basis
In this subsection, we focus on constructing a canonical basis based on the pre-canonical basis in App. B.1 under the maximal cut then towards the fully $\varepsilon$-factorized basis of the family. Similar to the Legendre family, we consider both of the two methods, with derivative basis in App. B.2.1 and constant period matrix in App. B.2.2.
B.2.1 The construction with derivative basis
@@
$\varpi_{0}(k)\equiv\frac{2}{\pi}\int_{-1}^{1}\frac{{\rm d}t}{\sqrt{P_{J}(t)}}=\frac{4}{\pi}K(k^{2})\,,\quad\varpi_{1}(k)\equiv\frac{2i}{\pi}\int_{1}^{\frac{1}{k}}\frac{{\rm d}t}{\sqrt{P_{J}(t)}}=\frac{2}{\pi}K(1-k^{2})\,,$ (117)
where $\varpi_{0}(k)$ and $\varpi_{1}(k)$ are holomorphic around degenerate points $k=0$ and $k=1$, respectively. The modular $\tau$ is defined as
$\tau\equiv\frac{i\varpi_{1}(k)}{\varpi_{0}(k)}=\frac{iK(1-k^{2})}{2K(k^{2})}\in{\mathbb{H}}\,.$ (118)
The corresponding derivative basis integrals are
$\tilde{I}_{1}=\frac{1}{\varpi_{0}(k)}I_{1}\,,\quad\tilde{I}_{n-1}=\frac{1}{\varepsilon}\frac{\varpi_{0}(k)^{2}}{\pi W_{k}}\partial_{k}\tilde{I}_{1}\,,$ (119)
@@
$W_{k}=\varpi_{0}\partial_{k}\varpi_{1}-\varpi_{1}\partial_{k}\varpi_{0}=-\frac{4}{\pi k(1-k^{2})}\,.$ (120)
The non-trivial elements in the transformation from pre-canonical basis $\vec{I}$ to $\tilde{\vec{I}}$ as ${\cal T_{\rm der}}$ are
$({\cal T}_{\rm der})_{1,1}$ $=\frac{1}{\varpi_{0}(k)}\,,$ (121a)
$({\cal T}_{\rm der})_{n-1,n-1}$ $=-\frac{1+\beta_{n+1}\varepsilon}{4\varepsilon}\varpi_{0}(k)\,,$ (121b)
$({\cal T}_{\rm der})_{n-1,j}$ $=-\frac{1}{4}\beta_{j+3}\frac{a_{j+3}(a_{j+3}^{2}-1)k^{2}}{\sqrt{P_{J}(a_{j+3})}}\varpi_{0}(k)\,,$ (121c)
$({\cal T}_{\rm der})_{n-1,n-2}$ $=-\frac{1}{4}k\varpi_{0}(k)\sum_{i=1}^{n}\beta_{i}a_{i}\,,$ (121d)
$({\cal T}_{\rm der})_{n-1,1}$ $=\frac{1}{4\varepsilon}\left\{(1-k^{2})\varpi_{0}(k)+k(1-k^{2})\varpi_{0}^{\prime}(k)-\varepsilon\sum_{i=1}^{n}\beta_{i}[1+k^{2}(a_{i}^{2}-1)]\varpi_{0}(k)\right\}\,,$ (121e)
where $32\leq j\leq n-$ 3 and other diagonal elements are equal to $1$. The non-zero elements in the coefficient matrix of ${\rm d}a_{i}$ in $\tilde{\bm{A}}^{(0)}$ are given by
$(\tilde{\bm{A}}^{(0)}_{a_{i}})_{i-3,1}$ $=-\frac{k[k(a_{i}^{2}-1)\varpi_{0}(k)+(1-k^{2})\varpi_{0}^{\prime}(k)]}{\sqrt{P_{J}(a_{i})}}\,,$ (122a)
$(\tilde{\bm{A}}^{(0)}_{a_{i}})_{n-1,1}$ $=\beta_{i}\frac{a_{i}k^{2}(1-k^{2})[\varpi_{0}(k)+k\varpi^{\prime}_{0}(k)]\varpi_{0}(k)}{4(a_{i}^{2}k^{2}-1)}\,,$ (122b)
$(\tilde{\bm{A}}^{(0)}_{a_{i}})_{n-1,i-3}$ $=\beta_{i}\frac{k(1-k^{2})[a_{i}^{2}k\varpi_{0}(k)+(a_{i}^{2}k^{2}-1)\varpi^{\prime}_{0}(k)]}{4(a_{i}^{2}k^{2}-1)\sqrt{P_{J}(a_{i})}}\,.$ (122c)
Elements $(\tilde{\bm{A}}^{(0)}_{a_{i}})_{i-3,1}$ can be removed with function $\vartheta_{0}(a_{i},k)$, and its explicit form is given by
$\vartheta_{0}(a_{i},k)\equiv-\frac{4\sqrt{P_{J}(a_{i})}}{\pi a_{i}}\Pi(1/a_{i}^{2},k^{2})\,,$ (123)
and its counterpart $\vartheta_{1}(a_{i},k)$ is
$\vartheta_{1}(a_{i},k)\equiv-\frac{2\sqrt{P_{J}(a_{i})}}{\pi a_{i}}\left[K(1-k^{2})+\frac{1}{a_{i}^{2}-1}\Pi\left(\frac{a_{i}^{2}(1-k^{2})}{a_{i}^{2}-1},1-k^{2}\right)\right]\,.$ (124)
The non-trivial elements in the transformation ${\cal T}$ from the pre-canonical basis $\tilde{\vec{I}}$ are
${\cal T}_{j,1}$ $=-\vartheta_{0}(a_{j+3},k)\,,$ (125a)
${\cal T}_{n-1,j}$ $=\frac{1}{4}\beta_{j+3}\left[\vartheta_{0}(a_{j+3},k)+\frac{a_{j+3}(a_{j+3}^{2}-1)k^{2}\varpi_{0}(k)}{\sqrt{P_{J}(a_{j+3})}}\right]\,,$ (125b)
${\cal T}_{n-1,1}$ $=-\frac{1}{8}\left\{\sum_{i=5}^{n}\beta_{i}\vartheta_{0}(a_{i},k)^{2}-\sum_{i=1}^{n}\beta_{i}[1+k^{2}(a_{i}^{2}-1)]\varpi_{0}(k)^{2}\right\}\,,$ (125c)
where diagonal elements are equal to $1$. With ${\cal T}_{\rm der}$ from (33) and ${\cal T}$ from (38), the transformation from pre-canonical basis $\vec{I}$ to canonical basis ${\cal T}_{\rm can}$ can be obtained directly
${\cal T}_{\rm can}={\cal T}\cdot{\cal T}_{\rm der}\,.$ (126)
The canonical basis and the corresponding connection matrix here share similar properties as the ones in the main text in Legendre family.
@@
${\cal T}_{\rm can}^{(\varepsilon)}=\begin{pmatrix}\frac{1}{\varpi_{0}(k)}&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ \frac{(1-k^{2})[\varpi_{0}(k)+k\varpi_{0}^{\prime}(k)]}{4\varepsilon}&0&\cdots&0&-\frac{(1+\beta_{n+1}\varepsilon)\varpi_{0}(k)}{4\varepsilon}\end{pmatrix}\,.$ (128)
B.2.2 The construction with constant period matrix

### v3

@@ Acknowledgements
We would like to thank Ming Lian for collaborations in the early stage of this work. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12375097, 12535003, 12347103, Science Foundation of China University of Petroleum, Beijing (No.2462025YJRC019), and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Construct $\varepsilon$-factorized basis with constant period matrix
In Sec. 2.3, we apply the method in Pogel:2022vat to construct the canonical basis. However, it is not the only method to find an $\varepsilon$-factorized basis. With a pre-canonical basis whose connection matrix in a special linear form, we can also try the strategy in vonManteuffel:2017hms ; Frellesvig:2021hkr .
The linear form as Eq. (22) is an essential step towards deriving an $\varepsilon$-factorized basis, which will be elaborated soon. In particular, $\bm{A}^{(0)}$ is closely related to the geometry behind the integral family and plays an important role in the later derivation. It is a good place here to put some remarks about it. For this purpose, we take $\varepsilon=0$ in this discussion. With a bit of abuse of notation, we are interested in the following equation:
${\rm d}\vec{I}^{{\rm pre},(0)}={\bm{A}^{(0)}}\,\vec{I}^{{\rm pre},(0)}\,.$ (111)
$\begin{aligned} {\rm d}\begin{pmatrix}I_{1}^{{\rm pre},(0)}\\ I_{n-1}^{{\rm pre},(0)}\end{pmatrix}=\begin{pmatrix}\frac{{\rm d}\lambda}{2(1-\lambda)}&-\frac{{\rm d}\lambda}{2\lambda(1-\lambda)}\\ \frac{{\rm d}\lambda}{2(1-\lambda)}&-\frac{{\rm d}\lambda}{2(1-\lambda)}\end{pmatrix}\begin{pmatrix}I_{1}^{{\rm pre},(0)}\\ I_{n-1}^{{\rm pre},(0)}\end{pmatrix}\end{aligned}\,$, (112)
which equivalently defines the elliptic curve while
${\rm d}I_{i-3}^{{\rm pre},(0)}=-\left[\frac{e_{i}(e_{i}-1)}{(1-\lambda)}\frac{{\rm d}\lambda}{2\sqrt{P_{L}(e_{i})}}+\frac{e_{i}{\rm d}e_{i}}{2\sqrt{P_{L}(e_{i})}}\right]I_{1}^{{\rm pre},(0)}+\left[\frac{e_{i}(e_{i}-1)}{\lambda(1-\lambda)}\frac{{\rm d}\lambda}{2\sqrt{P_{L}(e_{i})}}+\frac{{\rm d}e_{i}}{2\sqrt{P_{L}(e_{i})}}\right]I_{n-1}^{{\rm pre},(0)},$ (113)
with $5\leq i\leq n+1$, which is related to the marked points information in that curve. Denoting
$E=E(\sin^{-1}a_{i},\lambda)\,,\quad F=F(\sin^{-1}a_{i},\lambda)\,,\quad a_{i}=\sqrt{e_{i}/\lambda}$ (114)
${\rm d}I_{i-3}^{{\rm pre},(0)}$ $=-\left[{\rm d}(F-E)-E\frac{{\rm d}\lambda}{2(1-\lambda)}\right]I_{1}^{{\rm pre},(0)}+\left[{\rm d}F+\left(\frac{F}{2\lambda}-\frac{E}{2\lambda(1-\lambda)}\right){\rm d}\lambda\right]I_{n-1}^{{\rm pre},(0)}$ (115)
$=-{\rm d}\left[(F-E)I_{1}^{{\rm pre},(0)}-FI_{n-1}^{{\rm pre},(0)}\right],$
where the second equality in the above follows with the help of the elliptic information (112). This property motivates us to refine the master integrals $I_{i-3}^{\rm pre}$ for $5\leq i\leq n+1$, which are directly related to the marked points as
$I_{i-3}^{\rm pre}\to\big(\,\mathcal{T}_{1}\,\vec{I}^{\rm pre}\,\big)_{i-3}=I_{i-3}^{\rm pre}+\left[F(\sin^{-1}a_{i},\lambda)-E(\sin^{-1}a_{i},\lambda)\right]I_{1}^{\rm pre}-F(\sin^{-1}a_{i},\lambda)I_{n-1}^{\rm pre}\,.$ (116)
Now, the remaining task is to further transform $I_{1}^{\rm pre}$ and $I_{n-1}^{\rm pre}$, to rotate the corresponding $2\times 2$ block of $\bm{A}^{(0)}$ away, which is only $\lambda$ relevant. To do that, we only need to consider the corresponding $2\times 2$ block of the period matrix for $I_{1}^{\rm pre}$ and $I_{n-1}^{\rm pre}$, which is
@@
${\cal T}={\cal T}_{2}\,{\cal T}_{1}=\begin{pmatrix}E(1-\lambda)&0&\cdots&0&-K(1-\lambda)\\ F(\sin^{-1}a_{5},\lambda)-E(\sin^{-1}a_{5},\lambda)&1&\cdots&0&-F(\sin^{-1}a_{5},\lambda)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ F(\sin^{-1}a_{\infty},\lambda)-E(\sin^{-1}a_{\infty},\lambda)&0&\cdots&1&-F(\sin^{-1}a_{\infty},\lambda)\\ E(\lambda)-K(\lambda)&0&\cdots&0&K(\lambda)\end{pmatrix}\,$. (119)
It is straightforward to combine the transformation above with the pre-canonical basis in Eqs. (48) together with the Möbius transformation in Eq. (2.1), to obtain the corresponding $\varepsilon$-factorized basis at the integrand level. This supports the idea that our workflow can be combined with any algorithms to obtain integrand-level results, and then applied to generic integral families.
We’ve briefly discussed the asymptotic behaviors of the pre-canonical basis in the degenerate limit $c_{1}\to c_{2}$ and thus $\lambda\to 1$. It’s natural to consider the asymptotic behaviors of the $\varepsilon$-factorized basis as well. It suffices to study the asymptotic behaviors of the elliptic integrals appearing in the transformation matrices ${\cal T}_{1}$ and ${\cal T}_{2}$. We have the following:
@@
Note that $K(\lambda)$ and $F(\sin^{-1}a_{i},\lambda)$ are singular as $c_{1}\to c_{2}$, and they are logarithmically singular in this limit. These singularities do not cause problems for the basis integrals, since they cancel due to the fact that $\phi_{1}^{\rm pre}=\phi_{2}^{\rm pre}$ when $c_{1}\to c_{2}$; see Eq. (49). This cancellation also implies that the first $\varepsilon$-factorized basis integral vanishes in this limit. Similar to the pre-canonical basis, other $\varepsilon$-factorized basis integrals are ${\rm d}\log$-forms asymptotically in the neighborhood of the degenerate points $c_{1}=c_{2}$, as desired. From the demonstration above, we can see the construction in Sec. 2.3 is essentially different from the procedure here from two related perspectives. On the one hand, we introduce $\varepsilon$-dependence in the fiber transformations in the former construction, while the construction here does not. On the other hand, there are only complete elliptic integrals with a unique modulus arguments in the former case, while here we introduce incomplete elliptic integrals and both of complementary moduli arguments are needed. As a consequence, the construction here will introduce more elliptic integrals and undesirable asymptotic properties for the transformation. More relevant discussions can also be found in Frellesvig:2023iwr .
@@ B.1 The pre-canonical basis
$\phi_{1}^{\rm pre}$ $=\frac{{\rm d}t}{\sqrt{P_{J}(t)}}={\rm d}K(\sin^{-1}t,k^{2})\,$, (125a)
$\phi_{i-3}^{\rm pre}$ $=\frac{1}{t-a_{i}}\frac{\sqrt{P_{J}(a_{i})}}{\sqrt{P_{J}(t)}}{\rm d}t=\frac{1}{2}{\rm d}\log\left(\frac{1+\sqrt{\frac{(1-t^{2})(1-k^{2}a_{i}^{2})}{(1-a_{i}^{2})(1-k^{2}t^{2})}}}{1-\sqrt{\frac{(1-t^{2})(1-k^{2}a_{i}^{2})}{(1-a_{i}^{2})(1-k^{2}t^{2})}}}\right)-\frac{\sqrt{P_{J}(a_{i})}}{a_{i}}{\rm d}\Pi(1/a_{i}^{2},\sin^{-1}t,k^{2})\,$, (125b)
$\phi_{n-2}^{\rm pre}$ $=\frac{t}{\sqrt{P_{4J}(t)}}{\rm d}t=\frac{1}{2}{\rm d}\log\left(\frac{1+k\sqrt{\frac{1-t^{2}}{1-k^{2}t^{2}}}}{1-k\sqrt{\frac{1-t^{2}}{1-k^{2}t^{2}}}}\right)\,$, (125c)
$\phi_{n-1}^{\rm pre}$ $=\frac{1-k^{2}t^{2}}{\sqrt{P_{J}(t)}}{\rm d}t={\rm d}E(\sin^{-1}t,k^{2})\,$, (125d)
where $i=5,\cdots,n$ and the next to last basis differential $\phi_{n-2}^{\rm pre}$ is a ${\rm d}\log$-form. Then we can find that the differential equation is in a special linear form. The $\varepsilon^{0}$-part of the connection matrix with respect to $k$ is
@@
B.2 The transformation to the canonical ($\varepsilon$-factorized) basis
In this subsection, we focus on constructing the canonical basis based on the pre-canonical basis in App. B.1 under the maximal cut then towards the fully $\varepsilon$-factorized basis of the family. Similar to the Legendre family, we consider both of the two methods, with method Pogel:2022vat in App. B.2.1 and constant period matrix in App. B.2.2.
B.2.1 The construction of the canonical basis
Similarly, we decompose the transformation to the canonical basis into two parts,
${\cal T}_{\rm can}={\cal T}\,{\cal T}^{\prime}\,.$ (128)
@@ B.2.1 The construction of the canonical basis
$\varpi_{0}(k)\equiv\frac{2}{\pi}\int\limits_{-1}^{1}\frac{{\rm d}t}{\sqrt{P_{J}(t)}}=\frac{4}{\pi}K(k^{2})\,,\quad\varpi_{1}(k)\equiv\frac{2i}{\pi}\int\limits_{1}^{\frac{1}{k}}\frac{{\rm d}t}{\sqrt{P_{J}(t)}}=\frac{2}{\pi}K(1-k^{2})\,$, (129)
where $\varpi_{0}(k)$ and $\varpi_{1}(k)$ are holomorphic around degenerate points $k=0$ and $k=1$, respectively. Around degenerate point $k=1$, the modular $\tau$ is defined as
$\tau\equiv\frac{i\varpi_{0}(k)}{\varpi_{1}(k)}=\frac{2iK(k^{2})}{K(1-k^{2})}\in{\mathbb{H}}\,.$ (130)
The non-trivial elements in transformation $\cal T$$^{\prime}$ are
$({\cal T}^{\prime})_{1,1}$ $=\frac{1}{\varpi_{1}(k)}\,$, (131a)
$({\cal T}^{\prime})_{n-1,n-1}$ $=\frac{1+\beta_{n+1}\varepsilon}{4\varepsilon}\varpi_{1}(k)\,$, (131b)
$({\cal T}^{\prime})_{n-1,j}$ $=\frac{1}{4}\beta_{j+3}\frac{a_{j+3}(a_{j+3}^{2}-1)k^{2}}{\sqrt{P_{J}(a_{j+3})}}\varpi_{1}(k)\,,\quad({\cal T}^{\prime})_{j,j}=1\,,\quad(2\leq j\leq n-3)\,$, (131c)
$({\cal T}^{\prime})_{n-1,n-2}$ $=\frac{1}{4}k\varpi_{1}(k)\sum_{i=1}^{n}\beta_{i}a_{i}\,,\quad({\cal T}^{\prime})_{n-2,n-2}=1\,$, (131d)
$({\cal T}^{\prime})_{n-1,1}$ $=-\frac{1}{4\varepsilon}\left\{(1-k^{2})\varpi_{1}(k)+k(1-k^{2})\varpi_{1}^{\prime}(k)-\varepsilon\sum_{i=1}^{n}\beta_{i}[1+k^{2}(a_{i}^{2}-1)]\varpi_{1}(k)\right\}\,.$ (131e)
Under the transformation, the integral $I^{\rm pre}_{n-1}$ is transformed to
$\Big({\cal T}^{\prime}\,\vec{I}^{\rm pre}\Big)_{n-1}=\frac{1}{\varepsilon}\frac{\varpi_{1}(k)^{2}}{\pi W_{k}}\partial_{k}\bigg[\frac{1}{\varpi_{1}(k)}I_{1}\bigg]\,$, (132)
@@
$W_{k}=\varpi_{1}\partial_{k}\varpi_{0}-\varpi_{0}\partial_{k}\varpi_{1}=\frac{4}{\pi k(1-k^{2})}\,.$ (133)
The non-zero elements in $\tilde{\bm{A}}^{(0)}_{a_{i}}$ are given by
$(\tilde{\bm{A}}^{(0)}_{a_{i}})_{i-3,1}$ $=-\frac{k[k(a_{i}^{2}-1)\varpi_{1}(k)+(1-k^{2})\varpi_{1}^{\prime}(k)]}{\sqrt{P_{J}(a_{i})}}\,$, (134a)
$(\tilde{\bm{A}}^{(0)}_{a_{i}})_{n-1,1}$ $=-\beta_{i}\frac{a_{i}k^{2}(1-k^{2})[\varpi_{1}(k)+k\varpi^{\prime}_{1}(k)]\varpi_{1}(k)}{4(a_{i}^{2}k^{2}-1)}\,$, (134b)
$(\tilde{\bm{A}}^{(0)}_{a_{i}})_{n-1,i-3}$ $=-\beta_{i}\frac{k(1-k^{2})[a_{i}^{2}k\varpi_{1}(k)+(a_{i}^{2}k^{2}-1)\varpi^{\prime}_{1}(k)]}{4(a_{i}^{2}k^{2}-1)\sqrt{P_{J}(a_{i})}}\,$, (134c)
where $5\leq i\leq n$. Elements $(\tilde{\bm{A}}^{(0)}_{a_{i}})_{i-3,1}$ can be removed with function $\vartheta_{1}(a_{i},k)$, and its explicit form is given by
$\vartheta_{1}(a_{i},k)\equiv-\frac{2\sqrt{P_{J}(a_{i})}}{\pi a_{i}}\left[K(1-k^{2})+\frac{1}{a_{i}^{2}-1}\Pi\left(\frac{a_{i}^{2}(1-k^{2})}{a_{i}^{2}-1},1-k^{2}\right)\right]\,$, (135)
and its counterpart $\vartheta_{0}(a_{i},k)$ is
$\vartheta_{0}(a_{i},k)\equiv-\frac{4\sqrt{P_{J}(a_{i})}}{\pi a_{i}}\Pi(1/a_{i}^{2},k^{2})\,.$ (136)
The diagonal elements in $\cal T$} are all equal to $1$, and other non-trivial elements are
${\cal T}_{j,1}$ $=-\vartheta_{1}(a_{j+3},k)\,$, (137a)
${\cal T}_{n-1,j}$ $=-\frac{1}{4}\beta_{j+3}\left[\vartheta_{1}(a_{j+3},k)+\frac{a_{j+3}(a_{j+3}^{2}-1)k^{2}\varpi_{1}(k)}{\sqrt{P_{J}(a_{j+3})}}\right]\,,\quad(2\leq j\leq n-3)\,$, (137b)
${\cal T}_{n-1,1}$ $=\frac{1}{8}\left\{\sum_{i=5}^{n}\beta_{i}\vartheta_{1}(a_{i},k)^{2}-\sum_{i=1}^{n}\beta_{i}[1+k^{2}(a_{i}^{2}-1)]\varpi_{1}(k)^{2}\right\}\,.$ (137c)
With $\cal T$$^{\prime}$ from (131) and $\cal T$} from (137), the transformation from pre-canonical basis $\vec{I}^{\rm pre}$ to canonical basis ${\cal T}_{\rm can}$ can be obtained directly. Acting ${\cal T}_{\rm can}$ on Eqs. (125), we obtain a canonical basis in the Jacobi family. The canonical basis and the connection matrix here share similar properties as the ones in Legendre family.
@@
${\cal T}_{\rm can}^{(\varepsilon)}=\begin{pmatrix}\frac{1}{\varpi_{1}(k)}&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ -\frac{(1-k^{2})[\varpi_{1}(k)+k\varpi_{1}^{\prime}(k)]}{4\varepsilon}&0&\cdots&0&\frac{(1+\beta_{n+1}\varepsilon)\varpi_{1}(k)}{4\varepsilon}\end{pmatrix}\,$. (139)
B.2.2 The construction of an $\varepsilon$-factorized basis with constant period matrix